OpenStudy (anonymous):

Two boat landings are 2.7 km apart on the same bank of a stream that flows at 3.5 km/hr. A motorboat makes the round trip between the two landings in 2.9 hr. What is the speed of the boat relative to the water?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

ayee

OpenStudy (anonymous):

t1 + t2 = total time d/V1 =t 1 d/V2 =t2 d/V1 +d/V2 = total time 2.7 /(V-3.5) + 2.7/(V+3.5) = 2.9 \[ \frac{(V+3.5) + (V-3.5) }{ (V+3.5) (V-3.5) } = \frac{ 2.9 }{ 2.7 }\] \[{(V+3.5) + (V-3.5) } = \frac{ 2.9 }{ 2.7 } (V+3.5) (V-3.5)\] expand and group like terms. solve the quadratic for V